OpenStudy (anonymous):

MEDAL AND FAN!! Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (anonymous):

@kropot72 @KamiBug @Atschanz @ashhhrodriguez @ameenbriscoe13 @Crissy15 @brainbuzzed5646

OpenStudy (kropot72):

5x + y = 9 ............(1) 3x + 2y = 4 ..........(2) If you subtract 5x from both sides of equation (1), you can find y in terms of x as follows: 5x - 5x + y = y = 9 - 5x So we have: y = 9 - 5x ...........(3) Now you need to plug the expression for y in (3) into equation (2) and then solve to find the value of x.

OpenStudy (anonymous):

Anyone??

OpenStudy (anonymous):

the answer is the (-2,5)

OpenStudy (kropot72):

@knightmare6 Can you follow what I posted?

OpenStudy (anonymous):

Oh okay i got it

OpenStudy (kropot72):

So, plugging in the expression for y we get: 3x + 18 -10x = 4 .........(4) Can you solve (4) to find the value of x?

OpenStudy (anonymous):

I know the answer i need help on my other question i got (-2,5)

OpenStudy (anonymous):

what the there question

OpenStudy (kropot72):

That is not the correct choice.

OpenStudy (anonymous):

What do you mean then solve for (4) plug 4 in for x?

OpenStudy (kropot72):

3x + 18 -10x = 4 .........(4) You need to solve equation (4) to find the value of x. The first step in solving (4) is to combine like terms. 3x - 10x = -7x. So now we have: -7x + 18 = 4 ..............(5) Now you need to subtract 18 from both sides of equation (5). Can you do that?

OpenStudy (anonymous):

-7x=-14

OpenStudy (kropot72):

Good work! -7x=-14 .......(6) Now divide both sides of (6) by -7.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so its (2,-1)

OpenStudy (kropot72):

You have the correct choice. Well done!

OpenStudy (anonymous):

can you help me on my other question i tagged you in?

OpenStudy (kropot72):

Sorry, I must log out now.